Status Report

NASA Space Station On-Orbit Status 9 February 2006

By SpaceRef Editor
February 9, 2006
Filed under , , ,
NASA Space Station On-Orbit Status 9 February 2006

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2006) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.

Working jointly with ground engineers, the FE activated the Elektron oxygen generator, with the usual nitrogen (N2) purge of the BZh-8 Liquid Unit.  The electrolysis machine had been off since 2/2 for the EVA-15, the regeneration of the two BMP Harmful Impurities Removal system filter beds, and for using remaining Progress 19 stores.   [The cabin atmosphere was repressurized yesterday with air from Progress 20, exhausting 20P air reserves.  The repress raised total atmospheric pressure by 8.4 mmHg.]

In preparation for his second data collection session with the FOOT (Foot/Ground Reaction Forces during Space Flight) experiment, scheduled for tomorrow, CDR/SO McArthur assembled the equipment in the Lab, including preparations for the EMG (electromyography) calibration with camcorder/video recording, which will precede the day-long ops.

Wrapping up post-EVA closeout/cleanup activities, Tokarev and McArthur worked in the Service Module Transfer Compartment (SM PkhO) and DC1 Docking Compartment to restore all systems to their initial pre-EVA state.  This also included stowing Orlan-M suits and BSS control interface systems.

In addition, the FE completed post-EVA relocations of payload hardware from SM and FGB.   [This involved moving the Cryogem-03 cooler to the DC1 and re-activating it in thermostatic mode (at +20 degC) as well as relocating the BTKh-12 “Bioekologiya” kit, the “Biorisk-MSV” container, the ESA NOA (Nitric Oxide Analyzer) and Japan’s SCN (Space Cup Noodles) experiments, and “Matryoshka” hardware.]

Valery also collected and stowed EVA tools such as the KPU tool carrier, hammer, scissors, cutter, crow bar, wire ties, rubber restraints, etc.

Processing Status
Daily Mission
Return to Flight
Weekly Status
Weekly Science
Daily On-Orbit Status
Daily Crew Timeline
Soyuz | Progress
ISS News | ATV


By attaching jumper QDs (quick disconnects), Bill  re-established the ITCS LTL (Internal Thermal Control System/Lower Temperature Loop) connection of the EXPRESS rack 1(ER1) in support of the activation of the MAMS (Microgravity Acceleration Measuring System) for collecting structural vibration data during the station reboost next Saturday night (2/11).   [The ER1 LTL had been disconnected due to a slow leak of the coolant which was noted when ER1 was off.  ER1 activation supports the ongoing troubleshooting of the rack, and constraints are in place to minimize the amount of leakage into it.]

The CDR completed the monthly PEP (portable emergency provisions) safety inspection.  (Last time done: 1/10/06).   [The IMS (Inventory Management System)-supported inspection involves verification that PFEs (portable fire extinguishers), PBAs (portable breathing apparatus), QDMAs (quick-don mask assemblies) and EHTKs (extension hose/tee kits) are free of damage to ensure their functionality, and to track shelf life/life cycles on the hardware.  There are a total of 5 PBAs in the U.S. segment (USOS).  There is one EHTK, in the Lab.]

McArthur also conducted his weekly audit/inventory of the available CWCs (collapsible water containers) and their contents, to keep track of onboard water supplies.   [Updated “cue cards” based on Bill’s water calldowns are sent up every other week.  The current cue card lists two dozen water containers for the four types of water identified on board: technical water (for Elektron, flushing, hygiene; some CWCs found leaking), potable water (~180 liters), condensate water (for processing) and other (TCS fluid, EMU waste water).]

Tokarev serviced the BIO-5 Rasteniya-2 (“Plants-2”) experiment by rebooting and resetting its Lada-8 greenhouse control computer.

McArthur performed the daily routine maintenance of the SM’s environment control & life support system (SOZh) and later updated/edited the standard IMS (Inventory Management System) “delta file”, including locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

The CDR also conducted the daily atmospheric status check for ppO2 (Partial Pressure Oxygen) and ppCO2 (pp Carbon Dioxide), using the CSA-O2 (Compound Specific Analyzer-Oxygen sensor) and CDMK (CO2 Monitoring Kit).

In preparation for Saturday’s station reboost by the Progress 20 cargo ship, docked to the DC1 nadir port, Bill closed the Lab nadir window cover to protect it against thruster effluent.

After yesterday’s installation of the standard US-21 matching unit in the 20P, the hatches have been reopened.  Later today, TsUP-Moscow will conduct the standard dynamic (hot) testing of Russian thruster systems. [ISS attitude control will be handed over to the Russian MCS (motion control system) at 5:35pm EST.  There will be two test firings of Progress DPO (approach & attitude control) thrusters, each of 20 second duration: at 6:11pm using Progress DPO manifold #1 and at 7:43pm on manifold #2.  Control authority will be returned to the U.S. segment (USOS) at 8:00pm.]

The 20P test reboost on 2/11 from the DC1 nadir port is scheduled for 5:21pm EST on Daily Orbit 1, for a duration of 8 min 41.7 sec, resulting in a delta-V of 0.5 m/sec (1.64 ft/sec).   [USSTRATCOM has screened the post-maneuver trajectory against orbital debris data and has identified no post-maneuver close approaches.]

FE Tokarev conducted his daily check of the operation of the Japanese experiment GCF-JAXA (Granada Crystallization Facility) in the Russian TBU incubator, maintained at 20 degC, including a temperature check on its ART (automatic temperature recorder).   [This daily monitoring/temp checking, carried on the Russian voluntary “time available” task list, will continue until 4/30.]

The CDR filled out the regular weekly FFQ (food frequency questionnaire), his 13th, which keeps a log of his nutritional intake over time on special MEC (Medical Equipment Computer) software.   [On the MEC, Bill is using his personalized file that reflects the food flown for his Increment.  The FFQ records amounts consumed during the past week of such food items as beverages, cereals, grains, eggs, breads, snacks, sweets, fruit, beans, soup, vegetables, dairy, fish, meat, chicken, sauces & spreads, and vitamins.  IBMP/Moscow (Institute of Biomedical Problems, Russian: IMBP – Institute of Medico-Biological Problems) recommended average daily caloric value of the crew’s food intake is 2200-2300 cal.  If larger quantities of juices and fruits are taken into account, the value can go to 2400-2500 cal.]

As part of TVIS monthly maintenance, both crewmembers today conducted an in-depth inspection of their TVIS (Treadmill with Vibration Isolation & Stabilization) harnesses.   [Bill and Valery donned three harnesses (two of Expedition 12, one older from John Phillips) for checkout and photography for ground engineering analysis of any degradation and wear. The crew found no damage.  They were also provided with uplinked instructions for next week’s six-month TVIS maintenance, which includes roller bearing change-out and isolator wire rope replacement, currently timelined for 2/13-2/15.]

Later, the crew completed their regular 2.5-hr. physical exercise program on the TVIS, RED resistive exerciser and VELO bike with bungee cord load trainer.   [Valery’s daily protocol prescribes a strict four-day microcycle exercise with 1.5 hr on the treadmill in unmotorized mode and one hour on VELO plus load trainer (today: Day 3 of the first set).]

Afterwards, the CDR transferred the exercise data files to the MEC (Medical Equipment Computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

During yesterday’s EMU (Extravehicular Mobility Unit) activities in the U.S. Airlock, telemetry indicated that the secondary oxygen pack (SOP) pressure of EMU #3010 was slightly lower than expected (an apparent loss of 164 psi from 6238 psi measured last November).  A forward plan is being developed, but SOP pressure currently remains well above the Flight Rule minimum required for EVA (5410 psi).   [If the decrease is an actual leak (and not just a data bias caused by some other means), the allowable minimum would be reached on or about 7/20 this year, assuming constant current leak rate.  After that, the #3010 SOP could be swapped with the SOP from the third (spare) spacesuit #3013, leaving two EMUs usable for EVA by the station crew.]

McArthur’s activity yesterday to close out and clean up the PromISS-4 (Protein Crystal Growth Monitoring by Digital Holographic Microscope) experiment took longer than anticipated, so the final portion of this activity will be scheduled some time next week.  This includes the planned photography of the MSG (Microgravity Science Glovebox) multipurpose arm. 

Update on SuitSat (RadioSkaf):  A ground station yesterday reported receiving the signal from SuitSat, deployed by the crew during EVA-15 on 2/3/06.  The battery voltage is holding steady at approximately 26 Volts, and the radio has been operating for over 110 hours.

Today’s CEO (Crew Earth Observation) photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, were Sobat fans, SE Sudan (east of the White Nile River lie the Machar Marshes.  The wetlands are bounded on the east the Ethiopian Plateau from which much alluvial fan material originates. On this pass the crew attempted to map, from N to S, this transition area between wetlands and wooded, dissected hills), Internal waves, S Patagonian Shelf (on this pass the crew needed to look aft for sun glint enhancement of sea surface features, beginning with Bahia Blanca and continuing eastward to just north of the Falkland Islands), and Patagonian Glaciers (the crew had fair lighting this pass with clouds approaching from the SW.  They were to try for detailed views of the lower Southern Patagonian Ice Field, particularly the southernmost glaciers on the eastern flank).

To date, over 177,000 of CEO images have been taken in the first five years of the ISS.

  CEO photography can be viewed and studied at the websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 12 crew visit:

Expedition 12 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit  (as of this morning, 7:31am EST [= epoch]):

  • Mean altitude — 345.7 km
  • Apogee height — 351.5 km
  • Perigee height — 339.8 km
  • Period — 91.45 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0008692
  • Solar Beta Angle — 2.9 deg (magnitude increasing)
  • Orbits per 24-hr. day — 15.75
  • Mean altitude loss in last 24 hours — 75 m
  • Revolutions since FGB/Zarya launch (Nov. 98) — 41304

Events Ahead (all dates Eastern & tentative):

  • 02/11/06 — ISS test reboost (by 20P, @ 5:21pm EST, followed by maneuver to XPOP)
  • 02/22/06 — ISS reboost (by 19P; mvr. back to XPOP after burn)
  • 03/03/06 — Progress M-54/19P undocking & reentry (mvr. to LVLH XVV after undock)
  • 03/10/06 — ISS reboost (by SM thrusters; mvr. back to XPOP after burn)
  • 03/30/06 — Soyuz TMA-8/12S launch (Pavel Vinogradov/Russia, Jeffrey Williams/US, Marcos Pontes/Brazil)
  • 04/01/06 — Soyuz TMA-8/12S docking (DC1; mvr. to LVLH XVV after dock)
  • 04/09/06 — Soyuz TMA-7/11S undocking & reentry (mvr. to XPOP after undock)
  • 04/16/06 — Soyuz TMA-8/12S relocation (DC1 to FGB nadir port)
  • 04/24/06 — Progress M-56/21P launch
  • 04/26/06 — Progress M-56/21P docking
  • 05/03/06 — ULF1.1 launch (NET, not earlier than)
  • 06/15/06 — U.S. EVA (under review)
  • 06/19/06 — Progress M-55/20P undocking & reentry
  • 06/28/06 — Progress M-57/22P launch
  • 06/30/06 — Progress M-57/22P docking
  • 07/01/06 — 12A launch (under review)
  • 08/01/06 — Russian EVA-16 (under review)
  • 09/13/06 — Progress M-56/21P undocking & reentry
  • 09/14/06 — Soyuz TMA-9/13S launch
  • 09/16/06 — Soyuz TMA-9/13S docking (DC1)
  • 09/24/06 — Soyuz TMA-8/12S undocking & reentry
  • 10/18/06 — Progress M-58/23P launch
  • 10/20/06 — Progress M-58/23P docking
  • 12/19/06 — Progress M-57/22P undocking & reentry
  • 12/20/06 — Progress M-59/24P launch
  • 12/22/06 — Progress M-59/24P docking.

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at


SpaceRef staff editor.